# GraphPerf-RT: A Graph-Driven Performance Model for Hardware-Aware Scheduling of OpenMP Codes

### v2

@@ Abstract
Performance prediction for OpenMP workloads on heterogeneous embedded SoCs is challenging due to complex interactions between task DAG structure, control-flow irregularity, cache and branch behavior, and thermal dynamics; classical heuristics struggle under workload irregularity, tabular regressors discard structural information, and model-free RL risks overheating resource-constrained devices. We introduce GraphPerf-RT, the first surrogate that unifies task DAG topology, CFG-derived code semantics, and runtime context (per-core DVFS, thermal state, utilization) in a heterogeneous graph representation with typed edges encoding precedence, placement, and contention. Multi-task evidential heads predict makespan, energy, cache and branch misses, and utilization with calibrated uncertainty (Normal-Inverse-Gamma), enabling risk-aware scheduling that filters low-confidence rollouts. We validate GraphPerf-RT on three embedded ARM platforms (Jetson TX2, Jetson Orin NX, RUBIK Pi), achieving $R^{2}>0.95$ with well-calibrated uncertainty (ECE <0.05<0.05). To demonstrate end-to-end scheduling utility, we integrate the surrogate with four RL methods on Jetson TX2: single-agent model-free (SAMFRL), single-agent model-based (SAMBRL), multi-agent model-free (MAMFRL-D3QN), and multi-agent model-based (MAMBRL-D3QN). Experiments across 5 seeds (200 episodes each) show that MAMBRL-D3QN with GraphPerf-RT as the world model achieves 66% makespan reduction ($0.97\pm 0.35$s) and 82% energy reduction ($0.006\pm 0.005$J) compared to model-free baselines, demonstrating that accurate, uncertainty-aware surrogates enable effective model-based planning on thermally constrained embedded systems.
@@ 1 Introduction
Heterogeneous embedded Systems-on-Chip (SoCs) combine high-performance and energy-efficient cores with Dynamic Voltage and Frequency Scaling (DVFS). This architecture has created unprecedented opportunities for deploying parallel applications in resource-constrained environments such as autonomous vehicles, robotics, and edge computing. OpenMP, the dominant shared-memory parallel programming model, enables developers to express task-level parallelism through pragma-based annotations that the runtime system maps to available hardware resources. However, achieving optimal performance on these heterogeneous platforms requires careful scheduling decisions that balance execution time, energy efficiency, and thermal constraints. This presents a challenging optimization problem that current approaches fail to address comprehensively. Performance prediction for parallel workloads has traditionally relied on analytical models, simulation-based methods, and machine learning surrogates. Analytical models grounded in queueing theory and scheduling bounds such as Brent’s theorem provide theoretical insights but require simplifying assumptions that break down under irregular control flow and cache contention [13, 5]. Simulation-based approaches offer high fidelity but incur prohibitive runtime overhead for online scheduling decisions. Machine learning surrogates, including Graph Neural Networks (GNNs), can learn complex input-output mappings from data [39]. However, existing approaches such as ProGraML [10] and Ithemal [26] operate on static intermediate representations without incorporating runtime context, thermal state, or DVFS configurations that critically affect embedded system performance. Complementary approaches address related challenges through hierarchical multi-agent DVFS scheduling [31], zero-shot LLM-guided allocation [30], statistical feature-aware task allocation [32], and flow-augmented few-shot RL [33].
This paper addresses the problem of predicting multiple performance metrics for OpenMP task-parallel applications on heterogeneous embedded SoCs. These metrics include execution time, energy consumption, cache misses, branch misses, and CPU utilization under varying DVFS configurations, core allocations, and thermal conditions. The prediction model must provide accurate point estimates and calibrated uncertainty quantification to enable risk-aware scheduling decisions. This is essential for respecting thermal constraints and avoiding failures in safety-critical deployments. Furthermore, the model should generalize across benchmarks, input sizes, and hardware platforms without extensive retraining.
@@
This paper introduces GraphPerf-RT, a graph-grounded performance surrogate that addresses these challenges through a unified heterogeneous graph representation. This representation explicitly models the three-way interaction between application structure, hardware resources, and scheduling decisions. Task nodes encode CFG-derived code semantics extracted via the ARTIST2 Language for Flow analysis (ALF) and Execution Flow Graphs (EFGs). Resource nodes capture per-core DVFS state, utilization, and thermal headroom. Typed edges encode task-task precedence, task-resource placement, and resource-resource topology. A heterogeneous Graph Attention Network (GAT) architecture with type-specific encoders processes the graph through 3 to 6 attention layers. Evidential learning with Normal-Inverse-Gamma (NIG) distributions provides calibrated uncertainty, producing aleatoric and epistemic estimates in a single forward pass. The surrogate scores candidate configurations in batch, filters low-confidence proposals, and supports Dyna-style model-based planning where synthetic rollouts augment real-world samples. While this paper focuses on OpenMP on ARM-based embedded SoCs, the framework generalizes to other programming models. For CUDA and GPU workloads, kernels map to task nodes and Streaming Multiprocessors to resource nodes. For x86 systems, P-states replace ARM DVFS indices and Running Average Power Limit (RAPL) counters provide energy measurements. Comprehensive experiments on three embedded ARM platforms (NVIDIA Jetson TX2, Jetson Orin NX, and RUBIK Pi) span 42 benchmarks from BOTS and PolyBench. GraphPerf-RT achieves $R^{2}>0.95$ with well-calibrated uncertainty (Expected Calibration Error below 0.05), significantly outperforming tabular baselines and homogeneous GNN architectures. Integration with four RL methods (SAMFRL, SAMBRL, MAMFRL-D3QN, and MAMBRL-D3QN) demonstrates that MAMBRL-D3QN with GraphPerf-RT achieves 66% execution time reduction ($0.97\pm 0.35$s vs. $2.85\pm 1.66$s) and 82% energy reduction ($0.006\pm 0.005$J vs. $0.033\pm 0.026$J) compared to model-free baselines across 5 random seeds with 200 episodes each.
The main contributions of this paper are as follows:
@@
    A unified heterogeneous graph representation that jointly encodes OpenMP task DAG topology, CFG-derived code semantics, and runtime context (per-core DVFS, thermal state, utilization) through typed nodes and edges, enabling explicit modeling of cross-layer performance interactions.
@@
    An evidential multi-task prediction framework providing calibrated uncertainty quantification through Normal-Inverse-Gamma heads, producing both aleatoric and epistemic uncertainty in a single forward pass for risk-aware scheduling on embedded systems.
@@
    A practical scheduling integration demonstrating seamless combination with model-based RL, achieving 66% makespan and 82% energy improvements over model-free baselines while reducing on-device exploration through Dyna-style synthetic rollouts.
@@
    A reproducible evaluation framework including a complete data pipeline (OMPi + ALF-llvm + SWEET + telemetry), comprehensive experiments across three ARM platforms and 42 benchmarks, and statistical significance tests with 5-seed confidence intervals.
The remainder of this paper is organized as follows. Section 2 surveys related work on performance modeling, graph neural networks for systems, and uncertainty quantification. Section 3 formalizes the problem setup, defines the heterogeneous graph abstraction, and describes the data collection pipeline. Section 4 presents the GraphPerf-RT architecture including type-specific encoders, heterogeneous GAT layers, and evidential prediction heads. Section 5 reports experimental results on prediction accuracy, uncertainty calibration, and end-to-end RL scheduling evaluation. Section 6 concludes with limitations and future directions.
@@ 2 Related Work
Performance modeling for parallel systems spans analytical queueing/Petri-net models and simulators [11, 25], then statistical and neural surrogates [23, 17, 41, 7]. Black-box predictors (counters, instruction mixes) improve accuracy but lose structural signals and transfer. Learned models like Ithemal [26] use hierarchical LSTMs to predict x86 basic-block throughput from instruction sequences, achieving high accuracy on microarchitectural latency but operating on static assembly without capturing runtime parallelism, inter-task dependencies, or device/thermal state. Runtime autotuners (Apollo) adapt to input-dependent performance variations [34], yet remain kernel-focused. Our focus differs: we model OpenMP task DAGs on embedded SoCs, explicitly encoding dependencies, per-core DVFS, and thermal context to capture cross-layer effects needed for real-time decisions.
Graph neural networks enable learning on structured program and system graphs [42, 39]. ProGraML [10] constructs a unified graph representation from LLVM IR by combining control-flow, data-flow, and call edges, achieving strong results on compiler optimization tasks such as device mapping and thread coarsening; critically, it operates exclusively on static intermediate representation without runtime context, hardware state, or thermal conditions. Similarly, Ithemal [26] predicts microarchitectural latency from static instruction sequences but lacks parallelism structure, inter-task dependencies, or device state. Heterogeneous Graph Transformer (HGT) [16] extends graph attention to multiple node and edge types, providing a natural baseline for our heterogeneous task-resource graphs. However, HGT and similar architectures lack three critical capabilities for our embedded scheduling setting: (1) explicit edge decomposition that separates structural dependencies from execution features, (2) integration of thermal and DVFS context as part of the graph representation, and (3) evidential regression heads that produce calibrated uncertainty estimates for risk-aware decision-making. GNNs have also predicted hardware-dependent performance for ML models [38] and optimization effects [4]. Compiler cost models (Ansor, MetaSchedule, ROLLER) learn evaluators for tensor kernels and improve cross-device transfer [47, 37, 48, 45, 46], but operate below application-level task graphs and rarely encode runtime DVFS/thermal state. Recent work on transferable hardware embeddings includes Silhouette [29], which learns performance-conscious CPU embeddings for cross-platform prediction, while Glimpse [1] develops mathematical encodings of hardware specifications for neural compilation, provides an alternative to hand-crafted device sheets and could potentially enhance resource node features in our framework. However, these approaches focus on static kernel-level prediction without integrating dynamic per-core DVFS, thermal headroom, or application-level task DAG structure that we require for OpenMP scheduling under thermal constraints. In contrast, we construct a heterogeneous graph with task and resource nodes, typed edges with execution features, and CFG-informed encoders to improve zero-shot and few-shot generalization across benchmarks and boards while explicitly fusing runtime device/thermal context into the representation.
@@
Hardware-aware ML includes NAS with device constraints and cross-platform predictors [6], and performance modeling across CPU/heterogeneous configurations [40]. Energy/thermal control spans heuristic/optimization DVFS and RL [43]. Utilization-driven governors can mis-scale frequency by conflating compute and stalls [15, 24]. Model-free RL adapts but is sample- and heat-intensive on small boards [19]. Model-based control reduces sampling via learned dynamics and planning [27]. Recent approaches address these gaps through hierarchical multi-agent DVFS scheduling [31], statistical feature-aware task allocation [32], zero-shot LLM-guided allocation [30], and flow-augmented few-shot RL [33]. Our approach is complementary: a calibrated, graph-grounded evaluator supplies fast, hardware-aware rollouts to search/MARL under DVFS and thermal constraints.
@@
Uncertainty is critical for safe scheduling. Bayesian and ensemble methods provide confidence but add inference cost [22, 18]. Evidential learning yields calibrated intervals without sampling [36, 2]. We adopt evidential regression heads to expose confidence on makespan, energy, and counters and to gate synthetic rollouts, aligning with on-board, real-time planning.
@@
To our knowledge, GraphPerf-RT is the first unified graph representation that combines OpenMP task DAGs with runtime execution context (per-core DVFS states, thermal headroom, hardware counters) and employs evidential regression for calibrated uncertainty in safety-critical embedded scheduling. This integration enables risk-aware decisions under thermal constraints, critical for resource-constrained SoCs where overheating can trigger throttling or system failure, while maintaining the computational efficiency required for on-board deployment and real-time control.
@@ 3.1 Problem Formulation
[…]
Formally, we learn a mapping
$\mathcal{M}:\ (G,D,\mathcal{S},a)\ \mapsto\ (y,\ \mathcal{U}),$
where $\mathcal{U}$ contains the NIG distribution parameters $(\gamma,\nu,\alpha,\beta)$ for each target metric. Point estimates are the predictive means $\hat{y}_{k}=\gamma_{k}$, while prediction intervals derive from the NIG posterior. The uncertainty decomposition follows standard evidential learning formulations with aleatoric uncertainty given by $\beta_{k}/(\alpha_{k}-1)$ and epistemic uncertainty by $\beta_{k}/(\nu_{k}(\alpha_{k}-1))$.
The problem formulation assumes that task graphs are acyclic, the device sheet remains fixed during an episode, and thermal policies may dynamically shrink the feasible action set. Stochasticity from OS scheduling noise and co-runner interference induces variability in outcomes; the logged data supervises the model under these realistic conditions.
@@ 3.2 Heterogeneous Graph Abstraction
Device constants from the sheet $D$ are broadcast as shared features during node encoding. This includes DVFS tables, cache sizes, and governor flags that provide architectural context across all nodes. The explicit heterogeneous structure allows the GAT to learn interpretable attention patterns, attending strongly to thermal headroom when predicting energy or to critical path edges when predicting makespan.
3.3 Data Collection and Artifact Pipeline
Data preprocessing ensures clean, normalized features for model training. Duplicate tuples over $(G,\text{input},m,f)$ are removed to prevent data leakage. Outliers are filtered using median absolute deviation (MAD) on regression residuals. All features undergo z-score normalization computed per device, with global statistics retained separately for cross-device transfer experiments.
The dataset comprises 73,920 samples collected across three embedded platforms: 26,880 samples from RUBIK Pi with 8 ARM Cortex-A72 cores, 26,880 samples from Jetson Orin NX with 8 ARM Cortex-A78AE cores, and 20,160 samples from Jetson TX2 with 6 cores in a heterogeneous Denver plus Cortex-A57 configuration. Benchmarks span 42 programs from BOTS (alignment, fft, fib, floorplan, health, concom, knapsack, nqueens, sort, sparselu, strassen, uts) and PolyBench (30 kernels covering linear algebra, stencils, medley, and datamining categories). Each benchmark is executed in serial, tasks, and tied variants. Core masks sweep from 1 to 8 active cores on RUBIK Pi and Orin NX, and 1 to 6 on TX2. DVFS indices are evenly spaced across each device’s available frequency range.
Train, validation, and test splits follow a 60/20/20 ratio stratified by the tuple (benchmark, input size, core mask, DVFS index). This stratification ensures that test configurations are unseen during training, preventing the model from memorizing specific configuration outcomes. Notably, the same benchmark may appear across splits with different configurations, reflecting the realistic scenario where the model must generalize to new scheduling decisions for known workloads rather than predicting performance for entirely novel programs.
3.4 Learning Formulation and Outputs
For each target metric $k$, an evidential head outputs NIG parameters $(\gamma_{k},\nu_{k},\alpha_{k},\beta_{k})$. The predictive mean is $\hat{y}_{k}=\gamma_{k}$. Uncertainty decomposes into aleatoric and epistemic components:
$\text{Aleatoric}_{k}=\frac{\beta_{k}}{\alpha_{k}-1},\qquad\text{Epistemic}_{k}=\frac{\beta_{k}}{\nu_{k}(\alpha_{k}-1)}.$
Training minimizes the negative log marginal likelihood of the NIG distribution with evidence regularization to prevent overconfidence and a ranking term aligned to makespan-first scheduling objectives.
At runtime, the model processes batched candidate configurations to produce scores and prediction intervals. An uncertainty gate filters low-confidence proposals based on calibrated epistemic uncertainty thresholds. Selected actions execute on the device, producing a new telemetry row that optionally enters the replay buffer for continual learning. This inference pipeline integrates with the Dyna-style model-based RL methods evaluated in Section 5, where synthetic rollouts from the surrogate augment real-world samples to reduce on-device exploration.

### v3

# GraphPerf-RT: Graph-Driven Performance Modeling with Calibrated Uncertainty for OpenMP Scheduling on Heterogeneous Embedded SoCs

@@ Abstract
Autonomous AI agents on embedded platforms require real-time, risk-aware scheduling under resource and thermal constraints. Classical heuristics struggle with workload irregularity, tabular regressors discard structural information, and model-free reinforcement learning (RL) risks overheating. We introduce GraphPerf-RT, a graph neural network surrogate achieving deep learning accuracy at heuristic speeds (2–7ms). GraphPerf-RT is, to our knowledge, the first to unify task DAG topology, CFG-derived code semantics, and runtime context (per-core DVFS, thermal state, utilization) in a heterogeneous graph with typed edges encoding precedence, placement, and contention. Evidential regression with Normal-Inverse-Gamma priors provides calibrated uncertainty; we validate on makespan prediction for risk-aware scheduling. Experiments on three ARM platforms (Jetson TX2, Orin NX, RUBIK Pi) achieve $R^{2}=0.81$ on log-transformed makespan with Spearman $\rho=0.95$ and conservative uncertainty calibration (PICP $99.9\%$ at 95% confidence). Integration with four RL methods demonstrates that multi-agent model-based RL with GraphPerf-RT as the world model achieves 66% makespan reduction and 82% energy reduction versus model-free baselines, with zero thermal violations.
@@ 1 Introduction
Autonomous AI agents deployed on heterogeneous embedded Systems-on-Chip (SoCs) must make real-time scheduling decisions under stringent resource, energy, and thermal constraints. These agents, which power autonomous vehicles, robotic systems, and edge AI applications, require accurate performance prediction with calibrated uncertainty to ensure safe operation. Heterogeneous embedded SoCs combine high-performance and energy-efficient cores with Dynamic Voltage and Frequency Scaling (DVFS). This architecture has created unprecedented opportunities for deploying parallel applications in resource-constrained environments. OpenMP, the dominant shared-memory parallel programming model, enables developers to express task-level parallelism through pragma-based annotations that the runtime system maps to available hardware resources. However, achieving optimal performance on these heterogeneous platforms requires careful scheduling decisions that balance execution time, energy efficiency, and thermal constraints. This presents a challenging optimization problem that current approaches fail to address comprehensively. Performance prediction for parallel workloads has traditionally relied on analytical models, simulation-based methods, and machine learning surrogates. Analytical models grounded in queueing theory and scheduling bounds such as Brent’s theorem provide theoretical insights but require simplifying assumptions that break down under irregular control flow and cache contention [12, 5]. Simulation-based approaches offer high fidelity but incur prohibitive runtime overhead for online scheduling decisions. Machine learning surrogates, including Graph Neural Networks (GNNs), can learn complex input-output mappings from data [33]. However, existing approaches such as ProGraML [9] and Ithemal [24] operate on static intermediate representations without incorporating runtime context, thermal state, or DVFS configurations that significantly affect embedded system performance. Complementary approaches address related challenges through hierarchical multi-agent DVFS scheduling, zero-shot LLM-guided allocation, statistical feature-aware task allocation, and flow-augmented few-shot RL [Anonymous, 2026; details omitted for double-blind review].
This paper addresses the problem of predicting performance metrics for OpenMP task-parallel applications on heterogeneous embedded SoCs. The framework supports multiple targets including execution time, energy consumption, and hardware counters; we focus on makespan prediction under varying DVFS configurations, core allocations, and thermal conditions. The prediction model must provide accurate point estimates and calibrated uncertainty quantification to enable risk-aware scheduling decisions. This is essential for respecting thermal constraints and avoiding failures in safety-critical deployments. Furthermore, the model should generalize across benchmarks, input sizes, and hardware platforms without extensive retraining.
@@
Figure 1: GraphPerf-RT as a world model for safe, uncertainty-aware scheduling.
This paper introduces GraphPerf-RT, an AI technology that provides foundational infrastructure for safe, uncertainty-aware scheduling on embedded platforms. GraphPerf-RT addresses these challenges through a unified heterogeneous graph representation (see Fig. 1). Task nodes encode CFG-derived code semantics; resource nodes capture per-core DVFS state, utilization, and thermal headroom; typed edges encode task-task precedence, task-resource placement, and resource-resource topology. A heterogeneous Graph Attention Network (GAT) with type-specific encoders processes the graph through 3–6 attention layers. Evidential learning with Normal-Inverse-Gamma (NIG) distributions provides calibrated uncertainty in a single forward pass. The surrogate scores candidate configurations, filters low-confidence proposals, and supports Dyna-style model-based planning. Comprehensive experiments on three embedded ARM platforms (NVIDIA Jetson TX2, Jetson Orin NX, and RUBIK Pi) span 42 benchmarks from BOTS and PolyBench. GraphPerf-RT achieves $R^{2}=0.81$ on log-transformed makespan with Spearman $\rho=0.95$ for ranking and conservative uncertainty calibration (PICP $99.9\%$ at 95% confidence), significantly outperforming tabular baselines and homogeneous GNN architectures. Integration with four RL methods (SAMFRL, SAMBRL, MAMFRL-D3QN, and MAMBRL-D3QN) demonstrates that MAMBRL-D3QN with GraphPerf-RT achieves 66% execution time reduction ($0.97\pm 0.35$s vs. $2.85\pm 1.66$s) and 82% energy reduction ($0.006\pm 0.005$J vs. $0.033\pm 0.026$J) compared to model-free baselines across 5 random seeds with 200 episodes each.
The main contributions of this paper, which establishes GraphPerf-RT as AI infrastructure for autonomous embedded agents, are:
@@
    A unified heterogeneous graph representation for AI agents that jointly encodes OpenMP task DAG topology, CFG-derived code semantics, and runtime context (per-core DVFS, thermal state, utilization) through typed nodes and edges, enabling explicit modeling of cross-layer performance interactions critical for autonomous decision-making.
@@
    An evidential prediction framework supporting multi-task learning with calibrated uncertainty quantification through Normal-Inverse-Gamma heads, producing both aleatoric and epistemic uncertainty in a single forward pass. We validate on makespan prediction; the architecture extends to energy and hardware counter targets. This provides essential AI safety infrastructure for risk-aware scheduling on embedded systems where overconfident predictions can cause thermal violations or hardware damage.
@@
    A practical AI agent integration demonstrating seamless combination with model-based RL, achieving 66% makespan and 82% energy improvements over model-free baselines while reducing hazardous on-device exploration through Dyna-style synthetic rollouts enabled by GraphPerf-RT as a world model.
@@
    A reproducible AI4Tech evaluation framework including a complete data pipeline (OMPi + ALF-llvm + SWEET + telemetry), extensive experiments across three ARM platforms and 42 benchmarks, and statistical significance tests with 5-seed confidence intervals.
@@ 2 Related Work
Performance modeling spans analytical queueing/Petri-net models [10, 23], statistical surrogates [21, 15, 35, 7], and neural predictors. Black-box predictors using counters or instruction mixes improve accuracy but lose structural signals. Ithemal [24] predicts x86 basic-block throughput via hierarchical LSTMs but operates on static assembly without runtime parallelism or thermal state. Runtime autotuners like Apollo [28] adapt to input-dependent variations yet remain kernel-focused. We model OpenMP task DAGs on embedded SoCs, encoding dependencies, per-core DVFS, and thermal context for real-time decisions.
Graph Neural Networks.
GNNs enable learning on structured program graphs [37, 33]. ProGraML [9] builds unified graphs from LLVM IR for compiler tasks but operates on static IR without runtime or thermal context. HGT [14] extends graph attention to multiple node/edge types but lacks: (1) edge decomposition separating structure from execution features, (2) thermal/DVFS integration, and (3) evidential heads for calibrated uncertainty. GNNs have predicted hardware-dependent performance [32] and optimization effects [4]. We construct heterogeneous task-resource graphs with CFG-informed encoders and runtime context fusion.
Compiler Cost Models.
Ansor, MetaSchedule, and ROLLER learn kernel-level evaluators with cross-device transfer [42, 31, 43, 40, 41], but operate below task graphs and rarely encode DVFS/thermal state. Silhouette [27] learns CPU embeddings for cross-platform prediction; Glimpse [1] develops mathematical hardware encodings for neural compilation. These focus on static kernel prediction without dynamic DVFS, thermal headroom, or task DAG structure. GraphPerf-RT integrates typed edges, execution features, and runtime device/thermal context into a unified graph representation.
Hardware-Aware ML and DVFS.
@@
Hardware-aware ML includes NAS with device constraints [6] and cross-platform predictors [34]. Energy/thermal DVFS spans heuristics, optimization, and RL [38]. Utilization-driven governors mis-scale frequency by conflating compute and stalls [13, 22]. Model-free RL adapts but is sample- and heat-intensive [17]; model-based control reduces sampling via learned dynamics [25]. Recent work includes hierarchical multi-agent DVFS, feature-aware allocation, zero-shot LLM guidance, and flow-augmented RL [Anonymous, 2026]. GraphPerf-RT is complementary: a calibrated graph-grounded evaluator supplying fast rollouts to MARL under thermal constraints.
Uncertainty Quantification.
@@
Uncertainty is critical for safe scheduling. Bayesian and ensemble methods add inference cost [20, 16]. Evidential learning yields calibrated intervals without sampling [30, 2]. Early formulations suffered from evidence contraction [36], which we address through non-saturating regularization. We adopt evidential heads to expose confidence on makespan/energy and gate synthetic rollouts for real-time planning.
@@
To our knowledge, GraphPerf-RT is the first unified graph representation combining OpenMP task DAGs with runtime context (per-core DVFS, thermal headroom, counters) and evidential regression for calibrated uncertainty in embedded scheduling. This enables risk-aware decisions under thermal constraints while maintaining computational efficiency for on-board deployment.
@@ 3.1 Problem Formulation
Appendix N.1
We predict multiple performance metrics for OpenMP task-parallel applications on heterogeneous embedded SoCs under varying DVFS configurations. Each application is a task DAG $G=(V,E)$ with tasks $v\in V$ characterized by CFG-derived features and precedence edges $e\in E$ (spawn, join, or data dependencies). The platform has $C$ cores with discrete frequency levels $f_{i}\in\mathcal{F}_{i}$. The device sheet $D$ captures immutable hardware constants (core types, cache specs, DVFS tables), while runtime state $\mathcal{S}$ includes per-core frequencies, utilization, and thermal readings.
A scheduling action $a=(m,f,p)$ specifies core mask $m\in\{0,1\}^{C}$, DVFS vector $f$, and optional priority $p$. Given $(G,D,\mathcal{S},a)$, we learn $\mathcal{M}:(G,D,\mathcal{S},a)\mapsto(y,\mathcal{U})$, where $y$ can include makespan, energy, or hardware counters depending on the target configuration. We focus on makespan prediction with calibrated uncertainty $\mathcal{U}$ via NIG parameters $(\gamma,\nu,\alpha,\beta)$.
@@ 3.2 Heterogeneous Graph Abstraction
Figure 2: Heterogeneous graph with task ($V_{T}$), resource ($V_{R}$), and memory ($V_{M}$) nodes; typed edges encode cross-layer interactions.
Appendix N.2
Performance emerges from interactions between application structure, hardware state, and scheduling decisions. As illustrated in Fig. 2, our heterogeneous graph has three node types: Task nodes $V_{T}$ (orange in figure) encode CFG features, DAG topology (depth, distance-to-sink, centrality), and runtime snapshots. Resource nodes $V_{R}$ (teal) encode per-core DVFS state, utilization, and thermal headroom. Memory nodes $V_{M}$ (purple) encode cache hierarchy characteristics.
Four typed edges capture performance-critical cross-layer interactions: $E_{TT}$ (task-task precedence with critical-path flags), $E_{TR}$ (task-resource placement with affinity), $E_{RR}$ (resource-resource sharing/contention), and $E_{RM}$ (resource-memory bandwidth constraints). Device constants from $D$ are broadcast during encoding.
3.3 Data Collection Pipeline
Appendix N.3
OpenMP sources are compiled through OMPi, with ALF-llvm emitting LLVM IR and ALF files. We use ALF not as a legacy artifact, but as a specialized intermediate representation that performs semantic lifting—explicitly exposing control-flow structure (loop bounds, recursion depths, branch patterns) that provides semantically dense features standard LLVM IR processing would miss. SWEET generates CFG/call graphs from this lifted representation, and post-processing maps entities to OpenMP tasks with topological encodings. This approach bridges static analysis and runtime performance modeling by capturing control-flow invariants critical for accurate prediction. Runtime logging captures one CSV row per execution with telemetry: timestamps, DVFS indices, measured frequencies, performance counters (cycles, instructions, cache/branch misses), energy, and thermal readings.
The dataset comprises 73,920 samples across three platforms: RUBIK Pi (8 Cortex-A72 cores), Jetson Orin NX (8 Cortex-A78AE cores), and Jetson TX2 (6 cores, Denver+A57). We evaluate 42 benchmarks from BOTS and PolyBench with 60/20/20 train/val/test splits stratified by (benchmark, input, core mask, DVFS index).
3.4 Learning Formulation
Appendix N.4
Type-specific MLPs encode nodes into a common embedding space, with separate parameter sets for each node type to handle their distinct feature schemas. A heterogeneous GAT (3–6 layers) performs message passing over all edge types $E_{TT}$, $E_{TR}$, $E_{RR}$, and $E_{RM}$. Hierarchical pooling aggregates node embeddings per type and concatenates them to form graph representation $\mathbf{h}_{G}$. For each metric $k$, evidential heads output NIG parameters $(\gamma_{k},\nu_{k},\alpha_{k},\beta_{k})$ with the predictive mean $\hat{y}_{k}=\gamma_{k}$ and uncertainty decomposition: $\text{Aleatoric}_{k}=\beta_{k}/(\alpha_{k}-1)$ and $\text{Epistemic}_{k}=\beta_{k}/(\nu_{k}(\alpha_{k}-1))$. Training minimizes the negative log marginal likelihood with evidence regularization to prevent overconfidence. At runtime, batched configurations are scored, filtered by uncertainty gates, and executed with outcomes logged for continual learning.